Evaluate whether it violated or justify WARP conditionality: A consumer makes the

following choices (when you develop the matrix, show the calculation process for the first row)

a. At prices (p 1 ,p 2 )=($4,$4) the choice was (x 1 ,x 2 ) = (20,2).

b. At (p 1 ,p 2 )=($4,$2) the choice was (x 1 ,x 2 ) = (10,10).

c. At (p 1 ,p 2 )=($2,$4) the choice was (x 1 ,x 2 ) = (10,8)


1
Expert's answer
2021-08-16T14:47:04-0400


Calculation of the first row; At the price ($4,$4) and choice (20,4) ; 20x4 + 4x4 = 16

At the price ($4,$4) and choice (10,10) ; 4x10 + 4x10= 80

At the price ($4,$4) and choice (10,8) ; 10x4 + 4x8 = 72

The numbers in bold and diagonal are the numbers representing the cost of purchasing the bundles

while the numbers other than these shows the amount that the consumer c=would have spent if he does not purchase these are the bundles 

 

When bundle A is purchased, B and C are affordable but when B is purchased bundle A is not affordable and also, when Bundle C is purchased, Bundle A is not affordable

In situation A, Bundle A is directly revealed preferred over B and C

In situation B, Bundle B is directly revealed preferred over C

in situation C, Bundle C is directly revealed preferred over B and A 

Thus, this data does not violate WARP


The data does not violate WARP